Annie is five years older than twice her brother Dan's age. Write an alegraic expression to represent the sum of their ages.
Answer by Maths68(1474)

Let
Age of Annie = a
Age of Dan = d
Twice of Dan' age = 2d
Annie is five years older than twice her brother Dan's age means if we add 5 to the twice of Dan'age both Annie and Dan will be of equal age
Annie = twice of Dan's age + five years
a = 2d + 5
Sum of their ages = Annie'age + Dan'age
Sum of their ages = a + d
Put the value of a in above equation
Sum of their ages = (2d+5)+d
Sum of their ages = 2d+5+d
Sum of their ages = 3d+5
Let
Dan's age = 2
Annie age = 2d + 5 = 2*2 + 5 = 4+5=9
Sum of their ages = Annie's age + Dan's age
Sum of their ages = 9 + 2 = 11

Now put the values in expression for the sum of their ages
Sum of their ages = 3d+5
Sum of their ages = 3(2)+5 = 6+5=11